OpenStudy (anonymous):

ind an equation for the nth term of the arithmetic sequence. a14 = -33, a15 = 9

Nnesha (nnesha):

what's the formula for an arithmetic sequence ?

OpenStudy (anonymous):

not sure

Nnesha (nnesha):

general form equation for an arithmetic sequence ? you don't know ?

OpenStudy (anonymous):

No

Nnesha (nnesha):

\[\large\rm a_n=a_1 +(n-1)d\]you need this equation for an arithmetic sequence where n= term that we need nth a_1 = first term d=common difference in order to write the equation we have to find first term and common difference

Nnesha (nnesha):

formula to find common difference \[\rm d= a_2- a_1\]subtract the term by one previous term

OpenStudy (anonymous):

then you get 42 ohh i get it now

OpenStudy (anonymous):

an = -579 + 42(n - 1) so this is the answer right

Nnesha (nnesha):

substitute a_!4 and a_15 term into the equation when n=14 a_14 =-33 same with 15th term and then solve for d

Nnesha (nnesha):

how did you ge t -579 for first term ?

OpenStudy (anonymous):

wait what?

Nnesha (nnesha):

write two equations substitute n=14 a_14 =-33

OpenStudy (anonymous):

Thats what i did

Nnesha (nnesha):

\[\large\rm a_n=a_1+(n-1)d\] a_14 means n=14 remember n represent the number of the term so \[\rm a_{14}=a_1+(14-1)d\] and a_14 =-33 so plug that in \[\rm -33=a_1+(14-1)d\]

Nnesha (nnesha):

then that's what i asked how did you get -579 for first term......

OpenStudy (anonymous):

Oh i thought you were sayiing i was wrong thats why i said wait what

Nnesha (nnesha):

i don't know i'm saying show ur work how did you get -579